OpenStudy (anonymous):

im suuppeerr confused with this can someone help me understand how to do this (for medal and maybe fan)

OpenStudy (anonymous):

A pair of equations is shown below: y = 6x + 9 y = 7x + 7 What is the solution to the pair of equations? (–2, 21) (2, 21) (–2, –21) (2, –21)

OpenStudy (anonymous):

b

OpenStudy (anonymous):

is correct answer

Mehek (mehek14):

first, substitute the first equation into the second one

OpenStudy (anonymous):

want explination

OpenStudy (anonymous):

how do i solve lol yea ok @Mehek14 then what

Mehek (mehek14):

\(6x + 9= 7x + 7 \) then, combine like terms \(7x-6x=x\\9-7=2\\x=2\)

Mehek (mehek14):

so you have \((2,?)\) we need to find the y value so plug 2 in 6x + 9 \(6*2+9\\6*2=12\\12+9=21\\(2,21)\)

OpenStudy (anonymous):

ohhh ok thanks lol id fan you but i already did but if i have another question like this can you work it out with me you explained it good i just have different kind of that concept hard to explain lol

Mehek (mehek14):

ok